can you please help me write a quadratic equation with the roots 0,3 (in standad form)

Answer by stanbon(75887)

can you please help me write a quadratic equation with the roots 0,3 (in standad form)
-----
f(x) = (x-0)(x-3)
f(x) = x(x-3)
f(x) = x^2-3x
==================
Cheers,
Stan H.
==================